# Cosmological Stasis from Field-Dependent Decay

### v1

@@ 1 Introduction
It is generally expected that physics beyond the Standard Model (BSM) could manifest in the early stages of the cosmological timeline. In turn, the universe’s expansion history may also be influenced by BSM effects. Recently, it has been noticed that many BSM theories can give rise to a new kind of epoch dubbed “stasis” Dienes:2021woi in which cosmological abundances of different energy components, such as vacuum energy, matter, and radiation, remain constant over time when the universe continues to expand. It has been demonstrated that such stasis epochs emerge naturally as a global attractor of cosmological evolution in various contexts involving large towers of unstable states. These towers of states are predicted by many BSM theories, including those involving extra spacetime dimensions, string theories, and strongly-coupled gauge theories. For example, sequential decays of a tower of massive particles can lead to a stasis between matter and radiation Dienes:2021woi ; Dienes:2023ziv ; Halverson:2024oir . Similarly, continuous evaporation of a population of primordial black holes (PBHs) with an extended mass function can also lead to a matter/radiation stasis wherein the matter component is comprised of the PBHs Barrow:1991dn ; Dienes:2022zgd . The matter or radiation component can also be in stasis with vacuum energy if a tower of overdamped scalar fields can decay consecutively into radiation Dienes:2023ziv , or go through a series of overdamped/underdamped transitions Dienes:2023ziv ; Dienes:2024wnu . It is even possible that a triple stasis simultaneously between vacuum-energy, matter, and radiation can also arise if a tower of dynamical scalars experiences successively a combination of underdamping transition and decays Dienes:2023ziv .
Phenomenologically, modifications to the standard $\Lambda$CDM expansion history may lead to various observable signatures or constraints (see e.g.Batell:2024dsi for a recent review). As a highly non-standard cosmological epoch, stasis has the potential to unlock an array of new possibilities. For example, it has been shown that incorporating a PBH-induced stasis epoch into the cosmological timeline can alter the predictions for the inflationary observables in any model of inflation and leave distinct imprints on the spectrum of the stochastic gravitational background from inflation Dienes:2022zgd . Additionally, a stasis involving vacuum energy can lead to accelerated cosmic expansion, suggesting that the stasis epoch itself has the potential to be the inflationary era Dienes:2024wnu . Indeed, if proven viable, this so-called stasis inflation may open up a wealth of new theoretical avenues.
@@
Given these examples, it is tempting to claim that evidence for stasis can provide a smoking-gun signal for BSM theories that predict large towers of states. However, the presence of a tower is not the only way to realize stasis. For example, it has recently been found that a matter/radiation stasis can also emerge due to annihilations of a single particle species in thermal equilibrium with itself Barber:2024vui .
@@
Up to this point, in stasis scenarios, the decay rate of matter fields were assumed to be constant Dienes:2021woi ; Dienes:2023ziv ; Dienes:2024wnu ; Halverson:2024oir ; Barber:2024izt . In this paper, we point out that one may relax this condition. Indeed, in many BSM scenarios, couplings for various early-universe processes can acquire non-trivial time dependence by allowing some form of field dependence. A simple example is to promote the coupling constant $g$ that governs relevant decay processes to a dynamical quantity $g\sim(\phi/\Lambda)^{x}$. Indeed, such couplings arise in various scenarios for BSM physics, including the Froggatt-Nielsen model for addressing the flavor puzzle Froggatt:1978nt , the cosmology of flavons (see, e.g.Baldes:2016gaf ; Lillard:2018zts ; Chen:2019wnk ; Borboruah:2024eal ), early QCD confinement Ipek:2018lhm , baryogenesis Braconi:2018gxo ; Croon:2019ugf ; Elahi:2020pxl , dark matter production Buchmuller:2003is ; Berger:2020maa ; Batell:2021ofv ; Heurtier:2021rko ; Cheek:2024fyc , and the generation of gravitational waves Borboruah:2024eal ; Blasi:2024vew . With the coupling being dynamical, the energy transfer rate due to decay depends on the evolution of the field value $\phi$, which in turn is governed by the form of the potential $V(\phi)$. In the context of supergravity, the scalar field $\phi$ can couple to energy densities of various components in the universe and thus develop a potential of the form $V\sim H^{2}\phi^{2}$ — a potential commonly referred to as the Hubble mass potential or the Hubble mass term Copeland:1994vg ; Dine:1995uk ; Lyth:2004nx ; Kawasaki:2011zi ; Kawasaki:2012qm ; Kawasaki:2012rs . We find that the field evolution under this type of potential can lead to a decay rate that is power-law in time and thus can potentially give rise to a stasis epoch.
@@ 3.1 Basic ideas
in which $\Gamma_{\chi}$ is the decay width of $\chi$, and the pump term $P^{(\rho)}_{M\gamma}=\Gamma_{\chi}\rho_{\chi}$. Converting the energy densities to the cosmological abundances, Eq. (7) becomes
@@
where $C_{\Gamma}$ is a constant in time. Its specific time dependence is therefore controlled by the time evolution of $\phi$. Immediately, we notice that if the equation of motion of $\phi$ admits a stable solution $\phi\sim t^{-1/(2x)}$, the pump term $P_{\chi\gamma}=\Gamma_{\chi}\Omega_{\chi}$ in Eq. (12) shall scale like $1/t$ and potentially lead to stasis. Indeed, this type of pump term corresponds to the $n=1$ case in the generalized pumps $P_{ij}^{(\rho)}=Z\rho_{i}^{n}$ described in Ref. Barber:2024izt where $(i,j)$ labels different energy components, $Z$ is a positive prefactor. As anticipated, the $n=1$ pump requires a $1/t$ scaling behavior in $Z$ to realize stasis, which is exactly the scaling relation that we expect the decay width $\Gamma_{\chi}$ to acquire from the field evolution. Indeed, as we shall see, this scaling behavior of $Z$ with $t$ arises in a highly non-trivial manner as a consequence of the evolution of an additional dynamical variable — i.e., the field value of $\phi$.
@@
One possibility is that the potential $V$ is simply a polynomial function of $\phi$, and thus the power $-1/(2x)-2$ implies that $V\sim\phi^{2+4x}$. However, since for any non-vanishing positive integer value of $x$ the power of $\phi$ in the potential $2+4x\geq 6$, the insertion of $\phi_{s}$ into the equation of motion will necessarily introduce a relation between the coefficient $C$ and the parameter $\overline{\kappa}$. Since $\overline{\kappa}$ is a function of the stasis abundances $\overline{\Omega}_{M}$ and $\overline{\Omega}_{\gamma}$, this means that even if a matter/radiation stasis can be achieved via this type of potential, the stasis values, $\overline{\Omega}_{M}$ and $\overline{\Omega}_{\gamma}$, are sensitive to the initial value of the special solution $\phi_{s}(t^{(0)})$ where $t^{(0)}$ denotes the initial time. Such sensitivity is, in principle, acceptable. For example, the scenario discussed in Ref. Dienes:2024wnu also exhibits a similar sensitivity, wherein the initial field values of the tower components are connected to the fundamental scales of the scenario. By contrast, in our context, since $\phi$ is assumed to be a species with negligible abundance, its field value does not parametrize the overall scale of our problem. Thus, justifying this sensitivity with a natural motivation is highly nontrivial. We, therefore, choose not to consider this type of potential in this paper.
On the other hand, if the potential $V$ is quadratic in $\phi$ but at the same time contains a coefficient that is proportional to $1/t^{2}$ during stasis, the insertion of $\phi_{s}$ into Eq. (17) will not introduce any relation between the initial condition and the stasis abundance as the coefficient $C$ can be canceled. It is then reasonable to expect that the potential of the form $V\sim\phi^{2}/t^{2}$ can produce a stasis epoch that is insensitive to initial conditions. We shall, therefore, focus on this well-motivated possibility.
@@
where $\xi=\sqrt{(\overline{\kappa}-1)^{2}-8\lambda_{t}}$. Obviously, for $\xi>0$, the second term always grows larger, and becomes dominant relative to the first term for sufficiently large $t$. Identifying the growing solution with $\phi_{s}\sim t^{-1/(2x)}$, it is straightforward to find that
@@
Inserting this relation back into Eq. (18), we find that
$\phi(t)=t^{-\frac{1}{2x}}\times\begin{cases}C_{1}t^{-(\overline{\kappa}-1-1/x)}+C_{2}&\text{if~{}~{}~{}}\overline{\kappa}-1-1/x\geq 0\\ C_{1}+C_{2}t^{-(\overline{\kappa}-1-1/x)}&\text{if~{}~{}~{}}\overline{\kappa}-1-1/x<0\end{cases}\,.$ (20)
Apparently, only the first line of the above equation is consistent with our assumption that the growing solution scales as $t^{-1/(2x)}$.
@@
for a universe with $\Omega_{\chi}+\Omega_{\gamma}=1$. First, inserting Eq. (21) to the condition for the first line of Eq. (20) leads to
$x\geq\frac{1}{\overline{\kappa}-1}=\frac{6}{2+\overline{\Omega}_{\chi}}-1\,.$ (22)
Therefore, for any $\overline{\Omega}_{\chi}<1$, the above inequality implies that $x>1$. It also implies that only $x\geq 2$ can be consistent with the full range $10<\overline{\Omega}_{\chi}<$ 1. Thus, $x=2$ can be seen as the minimum case in which $x$ takes the smallest possible integer value. Using Eq. (19), Eq. (22) also give rise to a constraint on $x$ and $\lambda_{t}$ given by
@@
The physical condition $10<\overline{\Omega}_{\chi}<$ 1 (or equivalently 3/$23/2<\overline{\kappa}<$ 2 in Eq. (19)) will then put another constraint on the combination of $\lambda_{t}$ and $x$.
In Fig. 1, we plot contours of $\overline{\Omega}_{\chi}$ from Eq. (24) for different values of $x$ and $\lambda_{t}$. The white area represents unphysical values of the $\overline{\Omega}_{\chi}$, i.e.$\overline{\Omega}_{\chi}<0$ or $\overline{\Omega}_{\chi}>1$. The gray area bordered by the red line represents the area excluded by Eq. (23). As expected, the region in which $21<x<$ 2 cannot accommodate the full range of $\overline{\Omega}_{\chi}$. However, as long as $x\geq 2$, any physical value of $\overline{\Omega}_{\chi}$ is unconstrained by Eq. (23).
@@ 3.2 Motivation for the scalar potential $V\sim\phi^{2}/t^{2}$
where $\lambda$ encodes the coupling of $\phi$ to different energy densities. In general, $\lambda$ can be either positive or negative. However, in our context, $\phi\sim t^{-1/(2x)}$ is a decreasing function of time during stasis. This type of behavior cannot be realized if $\phi$ is governed by a Hubble mass potential with a negative $\lambda$. Therefore, $\lambda$ is required to be positive. Additionally, $\lambda$ should remain constant throughout stasis. Its relation to $\lambda_{t}$ is then given by $\lambda=9\lambda_{t}/\overline{\kappa}^{2}$.
@@
  • the initial “velocity” of the scalar field $\phi$ vanishes, i.e.$\left.{d\phi}/{dt}\right|_{t^{(0)}}=0$;
@@
Notice that we have used the superscript “$(0)$” to denote the initial values. In the left panel of Fig. 2, we find that the evolution of $\Omega_{\chi}$ can in general exhibit three types of behaviors:
@@
Indeed, the red and blue curves illustrate scenarios where the solution $\Omega_{\chi}=\overline{\Omega}_{\chi}$ is stable and persists over time, though one converges asymptotically, while the other approaches with oscillations. By contrast, the green curve shows a case in which the $\Omega_{\chi}=\overline{\Omega}_{\chi}$ solution is an unstable solution. Interestingly, these three types of behavior correspond to distinct ranges of $\overline{\Omega}_{\chi}$, while the exact ranges depend on the value of $x$. For our chosen minimal integer value, $x=2$, these ranges are approximately: $10.88\lesssim\overline{\Omega}_{\chi}<$ 1, where the system exhibits asymptotic convergence towards stasis; $0.6\lesssim\overline{\Omega}_{\chi}\lesssim 0.88$, where the solution undergoes oscillatory convergence to a stasis value, and $0<\overline{\Omega}_{\chi}\lesssim 0.6$, where undamped oscillations indicate that stasis is never achieved. By increasing $x$, the region exhibiting the stasis phenomenon can be extended toward smaller values of $\overline{\Omega}_{\chi}$. This dependence is shown in the right panel of Fig. 2, where stasis with $\overline{\Omega}_{\chi}\lesssim 0.6$ can emerge by increasing $x$ from 2 to larger integer values.

### v2

@@ 1 Introduction
It is generally expected that physics beyond the Standard Model (BSM) could manifest in the early stages of the cosmological timeline. In turn, the universe’s expansion history may also be influenced by BSM effects. Recently, it has been demonstrated that many BSM theories can give rise to a new kind of epoch dubbed “stasis” Dienes:2021woi ; Dienes:2023ziv in which cosmological abundances of different energy components remain constant over time when the universe continues to expand. Stasis has been shown to arise in cosmologies involving many possible combinations of such energy components — components including, for example, matter, radiation, or vacuum energy Dienes:2021woi ; Dienes:2023ziv ; Barrow:1991dn ; Dienes:2022zgd ; Dienes:2024wnu ; Halverson:2024oir ; Barber:2024vui ; Barber:2024izt . For instance, sequential decays of a tower of massive particles can lead to a stasis between matter and radiation Dienes:2021woi ; Dienes:2023ziv ; Halverson:2024oir . Such towers of states are predicted by many BSM theories, including those involving extra spacetime dimensions, string theories and strongly-coupled gauge theories. It has also been shown that initial conditions appropriate for stasis can be set up using gravitational interactions Long:2025wjw . Similarly, continuous evaporation of a population of primordial black holes (PBHs) with an extended mass function can also lead to a stasis between matter and radiation wherein the matter component is comprised of the PBHs Barrow:1991dn ; Dienes:2022zgd . The matter or radiation component can also be in stasis with vacuum energy if a tower of overdamped scalar fields can transfer its energy consecutively into radiation Dienes:2023ziv , or go through a series of overdamped/underdamped transitions Dienes:2023ziv ; Dienes:2024wnu . It is even possible that a triple stasis simultaneously between vacuum-energy, matter, and radiation can also arise if a tower of dynamical scalars experiences successively a combination of underdamping transitions and decays Dienes:2023ziv . In all these scenarios, it has been proven that stasis epochs can emerge naturally as a global attractor of cosmological evolution. This makes the stasis phenomenon potentially unavoidable within a wide range of parameter space in these BSM cosmologies.
[…]
@@
In general, the phenomenological aspects listed above depend on the properties of the stasis epoch. Such properties, in turn, reflect the structure of the underlying theory. For example, as found in Refs. Dienes:2021woi ; Dienes:2022zgd ; Dienes:2023ziv ; Dienes:2024wnu , the abundance configuration during stasis and the duration of stasis are ultimately determined by the polynomial scaling relations for the masses, decay rates, and cosmological abundances across the tower, as well as the hierarchy of relevant fundamental scales. Recent research employing machine-learning techniques has also extended this analysis to scenarios in which the mass spectra of the states within the tower exhibit exponential growth Halverson:2024oir . Given these examples, it is tempting to claim that evidence for stasis can provide a smoking-gun signal for BSM theories that predict large towers of states. However, the presence of a tower is not the only way to realize stasis. For example, it has recently been found that a matter/radiation stasis can also emerge due to annihilations of a single particle species in thermal equilibrium with itself Barber:2024vui .
@@
Up to this point, in stasis scenarios, the decay rate of matter fields were assumed to be constant Dienes:2021woi ; Dienes:2023ziv ; Dienes:2024wnu ; Halverson:2024oir ; Barber:2024izt . In this paper, we point out that one may relax this condition. Indeed, in many BSM scenarios, couplings for various early-universe processes can acquire non-trivial time dependence by allowing some form of field dependence. A simple example is to promote the coupling constant $g$ that governs relevant decay processes to a dynamical quantity $g\sim(\phi/\Lambda)^{x}$. Indeed, such couplings arise in various scenarios for BSM physics, including the Froggatt-Nielsen model for addressing the flavor puzzle Froggatt:1978nt , the cosmology of flavons (see, e.g.Baldes:2016gaf ; Lillard:2018zts ; Chen:2019wnk ; Borboruah:2024eal ), early QCD confinement Ipek:2018lhm , baryogenesis Braconi:2018gxo ; Croon:2019ugf ; Elahi:2020pxl , dark matter production Buchmuller:2003is ; Berger:2020maa ; Batell:2021ofv ; Heurtier:2021rko ; Cheek:2022yof ; Cheek:2024fyc , and the generation of gravitational waves Borboruah:2024eal ; Blasi:2024vew . With the coupling being dynamical, the energy transfer rate due to decay depends on the evolution of the field value $\phi$, which in turn is governed by the form of the potential $V(\phi)$. In the context of supergravity, the scalar field $\phi$ can couple to energy densities of various components in the universe and thus develop a potential of the form $V\sim H^{2}\phi^{2}$ — a potential commonly referred to as the Hubble mass potential or the Hubble mass term Copeland:1994vg ; Dine:1995uk ; Lyth:2004nx ; Kawasaki:2011zi ; Kawasaki:2012qm ; Kawasaki:2012rs . We find that the field evolution under this type of potential can lead to a decay rate that is power-law in time and thus can potentially give rise to a stasis epoch.
@@ 3.1 Basic ideas
in which $\Gamma_{\chi}$ is the decay width of $\chi$, and thus the pump term $P^{(\rho)}_{M\gamma}=\Gamma_{\chi}\rho_{\chi}$. Note that, the two equations above are coupled not only through the decay terms but also through the Hubble parameter $H$ via the Friedmann equation. We shall also assume that the Hubble parameter is dominated by $\rho_{\chi}$ and $\rho_{\gamma}$. Converting the energy densities to the cosmological abundances, Eq. (7) becomes
@@
where $C_{\Gamma}$ is a constant in time. Its specific time dependence is therefore controlled by the time evolution of $\phi$. Immediately, we notice that if the equation of motion of $\phi$ admits a stable solution $\phi\sim t^{-1/(2x)}$, the pump term $P_{\chi\gamma}=\Gamma_{\chi}\Omega_{\chi}$ in Eq. (12) shall scale like $1/t$ and potentially lead to stasis. Indeed, this type of pump term corresponds to the $n=1$ case in the generalized pumps $P_{ij}^{(\rho)}=Z\rho_{i}^{n}$ described in Ref. Barber:2024izt where $(i,j)$ labels different energy components, and $Z$ is a positive prefactor. As anticipated, the $n=1$ pump requires a $1/t$ scaling behavior in $Z$ to realize stasis, which is exactly the scaling relation that we expect the decay width $\Gamma_{\chi}$ to acquire from the field evolution. Indeed, as we shall see, this scaling behavior of $Z$ with $t$ arises in a highly non-trivial manner as a consequence of the evolution of an additional dynamical variable — i.e., the field value of $\phi$.
@@
One possibility is that the potential $V$ is simply a polynomial function of $\phi$, and thus the power 1/(2x)2-1/(2x)-2 implies that $V\sim\phi^{2+4x}$. However, since for any non-vanishing positive integer value of $x$ the power of $\phi$ in the potential $2+4x\geq 6$, the insertion of $\phi_{s}=Ct^{-1/(2x)}$ into the equation of motion will necessarily introduce a relation between the coefficient $C$ and the parameter $\overline{\kappa}$. Since $\overline{\kappa}$ is a function of the stasis abundances $\overline{\Omega}_{M}$ and $\overline{\Omega}_{\gamma}$, this means that even if a matter/radiation stasis can be achieved via this type of potential, the stasis values, $\overline{\Omega}_{M}$ and $\overline{\Omega}_{\gamma}$, are sensitive to the initial value of the special solution $\phi_{s}(t^{(0)})$ where $t^{(0)}$ denotes the initial time. For simplicity, we choose not to consider this type of potential in this paper. [^2] We note that the scenario discussed in Ref. Dienes:2024wnu also exhibits a similar sensitivity wherein the initial field values of the scalar fields in a tower are connected directly to the total energy density of the universe before stasis ends. The initial field values therefore controls the damping effects via the Hubble friction term. By contrast, in our context, since $\phi$ is assumed to be a species with negligible abundance before the end of stasis, its initial field value, and thus its initial energy density, has negligible impact to the Hubble parameter.
On the other hand, if the potential $V$ is quadratic in $\phi$ but at the same time contains a coefficient that is proportional to $1/t^{2}$ during stasis, the insertion of $\phi_{s}$ into Eq. (17) will not introduce any relation between the initial field value of $\phi_{s}$ and the stasis abundance as the coefficient $C$ can be canceled. It is then reasonable to expect that the potential of the form $V\sim\phi^{2}/t^{2}$ can produce a stasis epoch that is insensitive to initial conditions. We shall, therefore, focus on this well-motivated possibility.
@@
where $\xi=\sqrt{(\overline{\kappa}-1)^{2}-8\lambda_{t}}$. Obviously, for $\xi>0$, the second term always grows larger, and becomes dominant relative to the first term for sufficiently large $t$. Identifying the growing solution with $\phi_{s}\sim t^{-1/(2x)}$, it is straightforward to find that $\xi=\overline{\kappa}-1-1/x$, from which we obtain that
@@
for a universe with $\Omega_{\chi}+\Omega_{\gamma}=1$. First, inserting Eq. (20) to the condition that $\xi=\overline{\kappa}-1-1/x>0$, we find that
$x>\frac{1}{\overline{\kappa}-1}=\frac{4-\overline{\Omega}_{\chi}}{2+\overline{\Omega}_{\chi}}\,.$ (21)
Since $3/2<\overline{\kappa}<2$ for any $\overline{\Omega}_{\chi}\in(0,1)$, the above inequality implies that, in order to have $\xi>0$, one needs $x>1$ at least. It also implies that only $x\geq 2$ can be consistent with the full range $0<\overline{\Omega}_{\chi}<1$. Thus, $x=2$ can be seen as the minimum case in which $x$ takes the smallest possible integer value. Using Eq. (19), Eq. (21) also give rise to a constraint on $x$ and $\lambda_{t}$ given by
@@
Physically, this is simply a condition for the existence of a growing solution (with $\xi>0$ in Eq. (18)) that scales like $\phi_{s}\sim t^{-1/(2x)}$.
@@
The physical condition $0<\overline{\Omega}_{\chi}<1$ (or equivalently $3/2<\overline{\kappa}<2$ in Eq. (19)) will then put another constraint on the combination of $\lambda_{t}$ and $x$.
[…]
@@ 3.2 Motivation for the scalar potential $V\sim\phi^{2}/t^{2}$
where $\lambda$ encodes the coupling of $\phi$ to different energy densities. In general, $\lambda$ can be either positive or negative. However, in our context, since we require that $\phi\sim t^{-1/(2x)}$ in order to realize stasis, this type of behavior cannot be realized if $\phi$ is governed by a Hubble mass potential with a negative $\lambda$. More importantly, since we shall assume that the Hubble mass term is the dominant contribution to the full scalar potential, a negative $\lambda$ makes the potential unbounded from below — it would quickly drive the field to infinity which leads to an anti-de Sitter universe. Therefore, we shall require $\lambda$ to be positive throughout this paper. Additionally, $\lambda$ should remain constant throughout stasis. Its relation to $\lambda_{t}$ is then given by $\lambda=9\lambda_{t}/\overline{\kappa}^{2}$.
@@
[…]
@@
Notice that we have used the superscript “(0)(0)” to denote the initial values. In the left panel of Fig. 2, we find that the evolution of $\Omega_{\chi}$ can in general exhibit three types of behaviors:
@@
Indeed, the red and blue curves illustrate scenarios where the solution $\Omega_{\chi}=\overline{\Omega}_{\chi}$ is stable and persists over time, though one converges asymptotically, while the other approaches with oscillations. By contrast, the green curve shows a case in which the $\Omega_{\chi}=\overline{\Omega}_{\chi}$ solution is an unstable solution. Interestingly, these three types of behavior correspond to distinct ranges of $\overline{\Omega}_{\chi}$, while the exact ranges depend on the value of $x$. For our chosen minimal integer value, $x=2$, these ranges are approximately: $0.88\lesssim\overline{\Omega}_{\chi}<1$, where the system exhibits asymptotic convergence towards stasis; $0.6\lesssim\overline{\Omega}_{\chi}\lesssim 0.88$, where the solution undergoes oscillatory convergence to a stasis value, and $0<\overline{\Omega}_{\chi}\lesssim 0.6$, where undamped oscillations indicate that stasis is never achieved. By increasing $x$, the region exhibiting the stasis phenomenon can be extended toward smaller values of $\overline{\Omega}_{\chi}$. This dependence is shown in the right panel of Fig. 2, where stasis with $\overline{\Omega}_{\chi}\lesssim 0.6$ can emerge by increasing $x$ from 2 to larger integer values.